Circular No. 7854

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

CBAT@CFA.HARVARD.EDU (science)

URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2002 F1

S. Nakano, Sumoto, Japan, reports the visual discovery by Syogo Utsunomiya of a comet in morning twilight. Additional observations listed below were provided in response to a request by the Central Bureau.

2002 UT  α_{2000}  δ_{2000}  m_{1} Observer
Mar. 18.844  21^{h}44.^{m}6  +5^{o}26′  10.0 Utsunomiya
19.844  21 48.3  +6 14  10.0  _{”}
20.18863 21 49 43.^{s}38 +6 31 09′′.3 10.4 Sanchez
20.18912 21 49 43.48 +6 31 10.8 12.0  _{”}
20.18959 21 49 43.58 +6 31 12.5  _{”}
20.19008 21 49 43.69 +6 31 13.9  _{”}
20.19064 21 49 43.82 +6 31 15.5  _{”}
20.19115 21 49 43.95 +6 31 17.0  _{”}
20.48624 21 50 53.36 +6 46 18.2 12.0 Stevens
20.49153 21 50 54.55 +6 46 36.9 11.0  _{”}
20.49527 21 50 55.49 +6 46 48.1 10.8  _{”}
20.49962 21 50 56.48 +6 47 02.3 11.2  _{”}
20.50660 21 50 58.23 +6 47 22.1 11.4  _{”}
20.51162 21 50 59.27 +6 47 38.9 11.2  _{”}

S. Utsunomiya (Minami Oguni-machi, Aso-gun, Kumamoto-ken, Japan). 25×150 binoculars; comet diffuse with weak condensation; coma diameter 1′ on Mar. 18 and 1′.5 on Mar. 19.

S. Sanchez and J. Rodriguez (Mallorca). 0.4-m f/2 Schmidt telescope +

CCD. Tail 90′ in p.a. 210^{o}; coma diameter 2′. B. Stevens (Las Cruces, NM). 0.30-m f/6.3 reflector + CCD. Tail at least

12′.5 long (and ∼ 14^{o} wide in p.a.) centered at p.a. 237^{o}.

COMETS C/1999 N5 AND C/1999 N6 (SOHO) Further to IAUC 7853, reduced observations and orbital elements ap-

pear on MPEC 2002-F03 for two additional comets found by R. Kracht on the SOHO website. Both objects were visible in the C2 coronagraph; C/1999 N5, a member of the Marsden group (cf. IAUC 7832), was also visible in C3 images. C/1999 N6 may be associated with C/1999 M3.

Comet  1999 UT  α_{2000}  δ_{2000}
C/1999 N5 July 10.938  7^{h}12.^{m}8  +22^{o}10′
C/1999 N6  12.121  7 17.9  +22 33
2002 March 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green